169 g FeCr2O4, 298 g K2CO3 and an excess of O2 (g) are sealed in a reaction vessel and allowed to react at high temperature. The amount of K2CrO4 obtained is 194 g. Calculate the percent yield of K2CrO4.


4 FeCr2O4  +  8 K2CO3  + 7 O2 → 8 K2CrO4 + 2 Fe2O3 + 8 CO2 


( Molar mass: FeCr2O4 = 223.84, K2CO3=138.21, K2CrO4 =194.19 g/mole )


1
Expert's answer
2021-09-12T23:17:13-0400

Solution:

Calculate moles of each reactant:

(169 g FeCr2O4) × (1 mol FeCr2O4 / 223.84 g FeCr2O4) = 0.755 mol FeCr2O4

(298 g K2CO3) × (1 mol K2CO3 / 138.21 g K2CO3) = 2.156 mol K2CO3


Balanced chemical equation:

4FeCr2O4 + 8K2CO3 + 7O2 → 8K2CrO4 + 2Fe2O3 + 8CO2 

According to stoichiometry:

4 mol of FeCr2O4 reacts with 8 mol of K2CO3

Thus, 0.755 mol FeCr2O4 reacts with:

(0.755 mol FeCr2O4) × (8 mol K2CO3 / 4 mol FeCr2O4) = 1.51 mol K2CO3

However, initially there is 2.156 mol of K2CO3 (according to the task).

Therefore, FeCr2O4 acts as limiting reagent and K2CO3 is excess reagent.


According to stoichiometry:

4 mol of FeCr2O4 produces 8 mol of K2CrO4

Thus, 0.755 mol FeCr2O4 produces:

(0.755 mol FeCr2O4) × (8 mol K2CrO4 / 4 mol FeCr2O4) = 1.51 mol K2CrO4


Calculate the theoretical mass of K2CrO4:

(1.51 mol K2CrO4) × (194.19 g K2CrO4 / 1 mol K2CrO4) = 293.227 g K2CrO4


%yield of K2CrO4 = (practical mass of K2CrO4 / theoretical mass of K2CrO4) × 100%

%yield of K2CrO4 = (194 g K2CrO4 / 293.227 g K2CrO4) × 100% = 66.16% = 66.2%

%yield of K2CrO4 = 66.2%


Answer: The percent yield of K2CrO4 is 66.2%.